Let’s consider following choices that a policy maker face between the policy platforms – (3marks) i. A – (0, 250) ; B (250, 0) ; C (250, 350) - Discuss and compare the policy choices that policy maker is facing. Can you advocate any one of them over the other? Give reasons in support of your answer. ii. X – (0, 500) ; Y (500, 0) ; Z (700, 0) – Discuss and compare the policy choices that policy maker is facing. Can you advocate any one of them over the other? Give reasons in support of your answer. iii. If the policy maker is utilitarian, then combining the policy options of part i. and ii., what should have been the chosen policy? Explain your answer.


1
Expert's answer
2021-10-07T09:22:44-0400

I. If policy maker chooses A, then he will get the utility = 250 with one platform. While he will get utility =0 from the other platform.

If he chooses B, then again he will be getting utility =250 from one platform. While from the other platform, his utility = 0. So option A and B is more or less the same.

But if he chooses option C, he will get utility = 250 from one platform and utility = 350 from the other platform. This option is certainly better.

As per the monotonic preferences, he must choose option C because more is better.

ii. In the 2nd case too, option X and option Y provides utility = 500 from single platform. If we plot the indifference curve, the points will lie on the same indifference curve.

If he chooses option Z, he will get the utility =700, which imply that the policy maker will reach the higher indifference curve. And as per rule, higher the indifference curve, higher is the satisfaction.

So, policy maker must choose option Z, because his satisfaction level will be higher compared to option X and Y.

iii. Combining options of part i and part ii

Option 1 (0 , 500 + 250) = (0, 750)

Option 2 (500 + 250, 0) = (750, 0)

Option 3 (250 + 700, 350 +0) = (950, 350)

Again it is clear option 3 lies on the higher indifference curve.

So, again the policy maker will choose option 3, because higher the indifference curve, high will be the satisfaction.